During the current year, East Corporation had 3 million shares of common stock outstanding. 3,700, 7% convertible bonds, each with $1,000 face value, were issued at face amount at the beginning of the year. East reported income before tax of $7 million and net income of $4.2 million for the year. Each bond is convertible into 10 shares of common stock. What is diluted EPS? (Round your answer to 2 decimal places. Do not round your intermediate calculations.)

rev: 04_09_2018_QC_CS-123854, 05_25_2018_QC_CS-127783

Multiple Choice

  • $1.40.

  • $1.44.

  • $1.38.

  • $1.43.
Answer #1

Basic earning per share = 4200000/3000000 = 1.40 per share

Income tax rate = 2.8/7 = 40%

Diluted earning per share = Adjusted net income/Adjusted shares = (4200000+155400)/(3000000+37000) = 1.43

So answer is a) $1.40